# GW/DT invariants and 5D BPS indices for strips from topological recursion

### v1

@@ Acknowledgment
Special thanks go to Kohei Iwaki who first brought this problem to our attention. We also would like to thank Nafiz Ishtiaque and Saebyeok Jeong for discussions about related matters. We thank the organisers of the trimester program“Higher Rank Geometric Structures” at the Institut Henri Poincaré, Paris, where the collaboration ensued. SB acknowledges the working conditions and hospitality provided by CERN and by the University of Geneva. The work of S.B. has been supported in part by the ERC-SyG project “Recursive and Exact New Quantum Theory” (ReNewQuantum), which received funding from the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation program, grant agreement No. 810573. A.H. was funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) – project ID 551478549 and by the Swiss National Science Foundation (SNSF) through the Ambizione project “TRuality: Topological Recursion, Duality and Applications” under the grant agreement PZ00P2 223297. O.M. was supported by the fundamental junior IUF grant G752IUFMAR.
@@ 2. Topological Recursion for the strip geometries
This section will give some basic background on Topological Recursion and recent developments within TR which make it possible to derive the free energy explicitly via TR. We want to highlight that the actual derivation will not be carried out by the recursive definition of TR, but rather via a trick coming from a duality, the so-called xxyy duality.
@@ 2.1. Background on Topological Recursion
TR is a recursive procedure which generates from the initial data of a spectral curve an infinite family of multidifferentials $\omega_{g,n}$ on $n$ copies of the spectral curve. The definition [EO07b] of $\omega_{g,n}$ is recursive in 2g+n22g+n-2 with $g,n\in\mathbb{Z}_{\geq 0}$ and has a very beautiful pictorial interpretation. We refer to [Eyn14] for a classical overview and [Bou24] for a more recent one that focuses on algebraic structures.
The spectral curve, the initial data, is a tuple $(\Sigma,x,y,B)$ with $\Sigma$ a Riemann surface with two functions $x,y:\Sigma\to\mathbb{C}$. $B$ is a symmetric bidifferential on $\Sigma^{2}$ with second order pole on the diagonal, no residue and normalized along the $A$-cycles for a choice of a basis. Another description of the spectral curve typically used in the literature is given in terms of an affine curve, the vanishing locus of a polynomial equation in $x,y$, that is
@@
for $g\geq 2$ with $d\Phi(q)=\omega_{0,1}(q)$. For $g=0,1$, the definition is more involved and will not play any role here, see [EO07b] for details. The definition of $F_{g}=\omega_{g,0}$ arises from a more general principle valid in TR going from $\omega_{g,n+1}$ to $\omega_{g,n}$ by
@@
From the definition (2.2), several properties follow more or less directly [EO07b]. For 2g+n2>02g+n-2>0, all $\omega_{g,n}$:
@@
However, it is definitely not invariant under all, since swapping $x$ and $y$ leaves the symplectic form $dx\wedge dy$ (up to an overall factor of -1) invariant, but generates a completely different family of multidifferentials, which we will denote by $\omega_{g,n}^{\vee}$. This family has for instance poles at the ramification points of $y$ (rather than $x$) which were assumed to be different. Originally, the free energy $F_{g}$ was conjectured to be invariant under the xxyy swap [EO07b, EO07a] but shortly later observed not to be true [BS12]. In [EO13], it was further investigated that the difference between $F_{g}$ and the dual $F_{g}^{\vee}$ depends on the singular points of the curve, which were understood partially for meromorphic $x,y$. But there is still some understanding missing, there.
Recently, explicit formulas between the two families $\omega_{g,n}$ of the spectral curve $(\Sigma,x,y,B)$ and $\omega_{g,n}^{\vee}$ of the curve $(\Sigma,y,x,B)$ were first conjectured [BCGF+21], then proven for $g=0$ [Hoc24a] and in general in [ABDB+22], see also [Hoc23a] for equivalent formulations.
@@
$\omega_{g,1}$ $=-\omega_{g,1}^{\vee}+d\Bigg{(}\frac{\omega_{g-1,2}^{\vee}+\frac{1}{2}\sum_{\begin{subarray}{c}g_{1}+g_{2}=g\\ g_{i}>0\end{subarray}}\omega_{g_{1},1}^{\vee}\,\omega_{g_{2},1}^{\vee}}{dx\,dy}\Bigg{)}+\sum_{m=2}^{3g-1}\big{(}d\frac{1}{dx}\big{)}^{m}\Omega_{g,m},$ (2.6)
which will be of particular use later. Here, $\Omega_{g,m}$ is a one-form obtained from graphs decorated by $\omega_{g^{\prime},n^{\prime}}^{\vee}$ and derivatives thereof. We refer to [Hoc23a, ABDB+22] for the explicit expression. Note that $d\frac{1}{dx}$ first divides a differential form by $dx$ and then takes the exterior derivative, thus turns a one-form into a one-form.
The formula (2.5) can actually be understood as a universal duality, independent of TR, but proven to hold for any spectral curve $(\Sigma,x,y,B)$ with meromorphic $x,y$.
Several further developments building upon the theory of xxyy duality were achieved [ABDB+25c, ABDB+24c, ABDB+24d, ABDB+24a, ABDB+25b, ABDB+24b, ABDB+25a, BCGS25, Hoc23b, Hoc24b, Hoc25, HS25].
Most importantly for this article (since we are dealing with meromorphic $dx$ and $dy$ rather than meromorphic $x,y$) is the fact that the explicit xxyy duality formula (2.5) suggests an extension of TR that properly considers logarithmic poles of $x,y$ (i.e., simple poles of $dx,dy$). This new definition is called Log-TR, suggested in [Hoc24b] and further refined in [ABDB+24c].
Under certain circumstances (for instance, if a generic framing is taken for the mirror curve of a Calabi-Yau threefold) Log-TR reduces to TR, which motivated the celebrated remodeling conjecture [BKMnP09], later proved in [EO15, FLZ20].
For fixed framing $f=0$, the remodeling conjecture does not hold; see, for instance, [BS12] for a discussion. However, this situation is rescued by the setting of Log-TR, which again arises by enforcing the xxyy duality. Note further that Log-TR is compatible with limits beyond the families studied in [BBC+23].
@@
Then, the multidifferentials of Log-TR are defined by $\omega_{0,1}=y\,dx$, $\omega_{0,2}=B$, and, for negative Euler characteristic $\chi=2-2g-n<0$, recursively by
@@
The xxyy duals, $\omega_{g,n}^{\vee}$, are defined by swapping $x$ and $y$. One then takes residues around $p_{i}^{\vee}\in{Ram}(y)$ and considers the dual log-vital singular points of $x$, denoted $a_{1}^{\vee},\ldots,a_{M^{\vee}}^{\vee}$. The two families $\omega_{g,n}$ for the spectral curve $(\Sigma,x,y,B)$ and $\omega_{g,n}^{\vee}$ for $(\Sigma,y,x,B)$, both defined using Log-TR, are related via the expression (2.5).
@@ 2.2. Deriving the Free Energy
In principle, the result for $F_{g}$ is known for the curve (1.1). One can deduce it via circumventing with the remodeling conjecture [BKMnP09] to Gromov-Witten theory. It can be written in terms of Hodge class integrals on the moduli space of complex curves. However, a direct computation via Topological Recursion was never succeeded in the literature which will be rectified in this article to also gain further insight.
@@
In the setting, when two families $\omega_{g,n}$ and $\omega_{g,n}^{\vee}$ are related by (2.5) and $F_{g}$ (and $F_{g}^{\vee}$) is defined by (2.3) (and its dual version), an explicit formula for the difference $F_{g}-F_{g}^{\vee}$ is found in [Hoc25, Thm. 1.1]. This formula involves all $\Omega_{g,m}$ of (2.6). Since $F_{g}^{\vee}=0$, we actually find a new formula for $F_{g}$ in terms of the dual correlators $\omega_{g^{\prime},n^{\prime}}^{\vee}$ instead of $\omega_{g,n}$. The new formula for the curve (1.1) reads
@@
$F_{g}=\frac{1}{2-2g}\sum_{p\in\{1,\frac{1}{\alpha_{j}},\frac{1}{\beta_{j}}\}}\operatorname{Res}\displaylimits_{q\to p}y\sum_{m=2}^{3g-1}\big{(}d\frac{1}{dx}\big{)}^{m-1}\Omega_{g,m}.$ (2.11)
In the next step, we observe that $\sum_{m=2}^{3g-1}\big{(}d\frac{1}{dx}\big{)}^{m-1}\Omega_{g,m}$ is an exact one-form (even for every $m$). The residue of an exact one-form vanishes at every of its poles, that gives
@@
We can subtract this from (2.11) multiplied by $\log(p)$ for any $p\in\{1,\frac{1}{\alpha_{j}},\frac{1}{\beta_{j}}\}$
@@
where the integration constant is irrelevant. On the RHS, we have several terms, where each of them, is not an exact one-form, but the whole RHS is still exact. However, due to the carefully chosen constant $y(p)$, we have that (y(q)y(p))(y(q)-y(p)) vanishes linearly in (2.12). Thus, the first term on the RHS of (2.13) does not contribute, also $\omega_{g-1,2}^{\vee}=0$ for $g>0$. Finally, we end up with
@@
Sending the Kähler parameters to zero just the first term survives with the prefactor $1+r+s$ corresponding to the Euler characteristic of the toric Calabi-Yau. Thus, this computation gives an independent proof for [BS12, Conjecture 3] for strip geometries.
@@
Interchanging the $\hbar$-sum and the sum over $n$, one can resum for any $n$ the sum over $g$ (this is a formal computation),
$\frac{(\frac{\beta_{i}}{\beta_{j}})^{n}}{n}\sum_{g=1}^{\infty}\hbar^{2g-2}\frac{n^{2g-2}B_{2g}}{2g(2g-2)!}=\frac{(\frac{\beta_{i}}{\beta_{j}})^{n}}{n(e^{\hbar n/2}-e^{-n\hbar/2})^{2}}.$
@@
where the indices $(i,j)$ range from 0 to $s$ for $\beta$’s and $(i,j)$ range from 1 to $r$ for $\alpha$’s. This is the result in one sector. A proper mathematical derivation would be via resurgence applying Borel summation techniques from [GK21] and the final result would depend on the integration contour [ASTT23].
@@
In this section, we compare the explicit results obtained from TR in (2.15) with the free energy of topological string theory (1.3) via the identification (1.8) of 5D BPS index. Our proposal amounts to computing the 5D BPS indices for strip geometries using the machinery of TR and the newly established xxyy duality [ABDB+22, Hoc24a, Hoc25] together with Log-TR [Hoc24b, ABDB+24c].
Starting from $Z_{\mathrm{top}}$, following [OSY11], one can compute the BPS partition function in the noncommutative chamber as
@@
for these geometries, where the BPS partition function was shown to agree with matrix model computations. More precisely, one should perform Borel resummation on the asymptotic series (2.15) in different sectors of the Borel plane following [GHN23, ASTT23]. This way, one lands in one of the DT chambers [Sze08, JM08, OSY11]. Writing $Z_{\mathrm{top}}$ as in (1.10), one can extract the DT invariants. In fact, these expressions are asymptotic in $\hbar$, but are convergent in $Q_{i}$. The convergence conditions on $Q_{i}$ determine the DT chamber, as was also observed in the computations of Nekrasov partition functions [NO06, NO16]. In our expressions (2.15) and (2.16), we keep the parameters $\alpha_{i}$ and $\beta_{i}$ quite general, without specifying the exact combinations $Q_{i}$’s, the complex structure parameters that correspond to them. The right identification of the parameters is made once we consider the convergence of such series in $Q_{i}$’s.
@@
There exist a few already established techniques for computing these 5D indices using other methods. One of them is the “exponential networks” [ESW17, BLR19, BLR21, BLR23, BRSW25]. Briefly, the idea is to study the 3D-5D coupled system that we introduced in §1. Then one analyzes the wall-crossing of the 3D-5D BPS states. The information of the 5D BPS spectrum can be extracted by constructing a nonabelianization map for the exponential networks. In other words, the wall-crossing formula for 3D-5D BPS states, which is of the universal Kontsevich-Soibelman form [KS08], one can extract 5D BPS indices, by looking at wall-crossing of some particular 3D-5D BPS states and this is achieved by analyzing the topologies of the exponential networks associated to the corresponding toric CY 3-folds. There are several other techniques to compute these invariants, for example using the BPS quiver and studying the moduli space of stable representations therein [BMP21, MP25]. Moreover, there exist techniques for computing them using supersymmetric localization [NO06, AK09] or using methods from matrix model [OSY11, Sul11], gluing of topological vertex [IKP06], to name a few. The strip geometries are rather well-studied and many of the techniques mentioned above can extract the 5D BPS indices quite efficiently. Each of these methods brings forth certain aspects of mirror symmetry, such as the counting in [BLR21, BLR20] are related to A-branes which are objects in Fukaya category, whereas the BPS quivers in [MP25] are related to counting of B-branes which are objects in the derived category of coherent sheaves. This exemplifies one aspect of homological mirror symmetry.
@@
Advent of supersymmetric localization [NO16, Nek17, JL25] gave rise to yet another set of powerful techniques to study these 5D indices. In this situation, one needs to study certain moduli space which are multiplicative Higgs bundles. This is a more direct path towards quantization and as we shall see in an upcoming work [BH] topological recursion is useful in that context too. Finally, there are several studies to understand the integrable structures of these 5D theories [Nek03, Nek04, OSY11].
@@
The direct computation using TR has also its own implications. GW invariants can be reformulated using “remodeling conjecture” [BKMnP09] on CY 3-fold $\mathcal{X}$. The result $F=\sum_{g\geq 0}\hbar^{2g-2}F_{g}$ is an asymptotic series in $\hbar$, where coefficients of each of the terms in $F_{g}$ in (2.15) encode the Gopakumar-Vafa invariants, for strip geometries. These GV invariants have clear interpretation in M-theory, as counting of M2-branes. The counts are also equal to the 5D BPS indices. In this work, we give a proof of these statements through topological recursion. As a check, we provide a few checks by comparing the free energy (2.15) and 5D invariants computed using other techniques, showing that they are in accordance with each other. One crucial simplification for the cases here is that, because there is no compact four cycle, the wall-crossing of the 5D BPS states is tractable and we can extract them by essentially analyzing one of the wall-crossing chambers.

### v2

@@ Acknowledgment
Special thanks go to Kohei Iwaki who first brought this problem to our attention. We also would like to thank Nafiz Ishtiaque and Saebyeok Jeong for discussions about related matters. We thank the organizers of the trimester program“Higher Rank Geometric Structures” at the Institut Henri Poincaré, Paris, where the collaboration ensued. SB acknowledges the working conditions and hospitality provided by CERN and by the University of Geneva. The work of S.B. has been supported in part by the ERC-SyG project “Recursive and Exact New Quantum Theory” (ReNewQuantum), which received funding from the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation program, grant agreement No. 810573. A.H. was funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) – project ID 551478549 and by the Swiss National Science Foundation (SNSF) through the Ambizione project “TRuality: Topological Recursion, Duality and Applications” under the grant agreement PZ00P2 223297. O.M. was supported by the fundamental junior IUF grant G752IUFMAR.
@@ 2. Topological Recursion for the strip geometries
This section will give some basic background on Topological Recursion and recent developments within TR which make it possible to derive the free energy explicitly via TR. We want to highlight that the actual derivation will not be carried out by the recursive definition of TR, but rather via a universal duality, the so-called xxyy duality, which enables us to compute the free energy efficiently.
@@ 2.1. Background on Topological Recursion
TR is a recursive procedure which generates from the initial data of a spectral curve an infinite family of multidifferentials $\omega_{g,n}$ on $n$ copies of the spectral curve. The definition [35] of $\omega_{g,n}$ is recursive in 2g+n22g+n-2 with $g,n\in\mathbb{Z}_{\geq 0}$ and has a very beautiful pictorial interpretation. We refer to [34] for a classical overview and [27] for a more recent one that focuses on algebraic structures.
The spectral curve, the initial data, is a tuple $(\Sigma,x,y,B)$ with $\Sigma$ a Riemann surface with two functions $x,y:\Sigma\to\mathbb{C}$. $B$ is a symmetric bidifferential on $\Sigma\times\Sigma$ with second order pole on the diagonal, no residue and normalized along the $A$-cycles for a choice of a basis. Another description of the spectral curve typically used in the literature is given in terms of an affine curve, the vanishing locus of a polynomial equation in $x,y$, that is
@@
for $g\geq 2$ with $d\Phi(q)=\omega_{0,1}(q)$. For $g=0,1$, the definition is more involved and will not play any role here, see [35] for details. The definition of $F_{g}=\omega_{g,0}$ arises from a more general principle valid in TR going from $\omega_{g,n+1}$ to $\omega_{g,n}$ by
@@
From the definition (2.2), several properties follow more or less directly [35]. For 2g+n2>02g+n-2>0, all $\omega_{g,n}$:
@@
However, it is definitely not invariant under all, since swapping $x$ and $y$ leaves the symplectic form $dx\wedge dy$ (up to an overall factor of -1) invariant, but generates a completely different family of multidifferentials, which we will denote by $\omega_{g,n}^{\vee}$. This family has for instance poles at the ramification points of $y$ (rather than $x$) which were assumed to be different. Originally, the free energy $F_{g}$ was conjectured to be invariant under the xxyy swap [35, 32] but shortly later counter-examples were found [26]. In [33], it was further investigated that the difference between $F_{g}$ and the dual $F_{g}^{\vee}$ depends on the singular points of the curve, which were understood partially for meromorphic $x,y$.
Recently, explicit formulas between the two families $\omega_{g,n}$ of the spectral curve $(\Sigma,x,y,B)$ and $\omega_{g,n}^{\vee}$ of the curve $(\Sigma,y,x,B)$ were first conjectured [23], then proven for $g=0$ [46] and in general in [3], see also [44] for equivalent formulations.
@@
$\omega_{g,1}$ $=-\omega_{g,1}^{\vee}+d\Bigg(\frac{\omega_{g-1,2}^{\vee}+\frac{1}{2}\sum_{\begin{subarray}{c}g_{1}+g_{2}=g\\ g_{i}>0\end{subarray}}\omega_{g_{1},1}^{\vee}\,\omega_{g_{2},1}^{\vee}}{dx\,dy}\Bigg)+\sum_{m=2}^{3g-1}\left(d\frac{1}{dx}\right)^{m}\Omega_{g,m},$ (2.6)
which will be of particular use later. Here, $\Omega_{g,m}$ is a one-form obtained from graphs decorated by $\omega_{g^{\prime},n^{\prime}}^{\vee}$ and derivatives thereof. We refer to [44, 3] for the explicit expression. Note that $d\frac{1}{dx}$ first divides a differential form by $dx$ and then takes the exterior derivative, thus turns a one-form into a one-form.
The formula (2.5) can actually be understood as a universal duality, independent of TR, but proven to hold for the TR-differentials $\omega_{g,n}$ for any spectral curve $(\Sigma,x,y,B)$ with meromorphic $x,y$.
Several further developments building upon the theory of xxyy duality were achieved [10, 6, 7, 4, 9, 5, 8, 24, 45, 47, 48, 43].
Most importantly for this article (since we are dealing with meromorphic $dx$ and $dy$ rather than meromorphic $x,y$) is the fact that the explicit xxyy duality formula (2.5) suggests an extension of TR that properly considers logarithmic poles of $x,y$ (i.e., simple poles of $dx,dy$). This new definition is called Log-TR, suggested in [47] and further refined in [6].
Under certain circumstances (for instance, if a generic framing is taken for the mirror curve of a Calabi-Yau threefold) Log-TR reduces to TR, which motivated the celebrated remodeling conjecture [25], later proved in [36, 38].
For fixed framing $f=0$ or f=1f=-1, the remodeling conjecture does not hold; see, for instance, [26] for a discussion. However, this situation is rescued by the setting of Log-TR, which again arises by enforcing the xxyy duality. Note further that Log-TR is compatible with some limits beyond the families studied in [22].
@@
Then, the multidifferentials of Log-TR are defined by $\omega_{0,1}=y\,dx$, $\omega_{0,2}=B$, and, for negative Euler characteristic $\chi=2-2g-(n+1)<0$, recursively by
@@
The xxyy duals, $\omega_{g,n}^{\vee}$, are defined by swapping $x$ and $y$. One then takes residues around $p_{i}^{\vee}\in{Ram}(y)$ and considers the dual log-vital singular points of $x$, denoted $a_{1}^{\vee},\ldots,a_{M^{\vee}}^{\vee}$. The two families $\omega_{g,n}$ for the spectral curve $(\Sigma,x,y,B)$ and $\omega_{g,n}^{\vee}$ for $(\Sigma,y,x,B)$, both defined using Log-TR, are related via the expression (2.5), see [6].
@@ 2.2. Deriving the Free Energy
In principle, the result for $F_{g}$ is known for the curve (1.1). One can deduce it via circumventing with the remodeling conjecture [25] to Gromov-Witten theory. It can be written in terms of Hodge class integrals on the moduli space of complex curves. Although the free energy $F_{g}$ was computed using the topological vertex method [50] a direct computation via Topological Recursion was never performed in the literature which we provide in this article and as a bonus we also gain further insights.
@@
In the setting, when two families $\omega_{g,n}$ and $\omega_{g,n}^{\vee}$ are related by (2.5) and $F_{g}$ (and $F_{g}^{\vee}$) is defined by (2.3) (and its dual version), an explicit formula for the difference $F_{g}-F_{g}^{\vee}$ is found in [48, Thm. 1.1]. This formula involves all $\Omega_{g,m}$ of (2.6). Since $F_{g}^{\vee}=0$, we actually find a new formula for $F_{g}$ in terms of the dual correlators $\omega_{g^{\prime},n^{\prime}}^{\vee}$ instead of $\omega_{g,n}$. The new formula for the curve (1.1) reads
@@
$F_{g}=\frac{1}{2-2g}\sum_{p\in\{1,\frac{1}{\alpha_{j}},\frac{1}{\beta_{j}}\}}\operatorname{Res}\displaylimits_{q\to p}y\sum_{m=2}^{3g-1}\left(d\frac{1}{dx}\right)^{m-1}\Omega_{g,m}.$ (2.11)
In the next step, we observe that $\sum_{m=2}^{3g-1}\big(d\frac{1}{dx}\big)^{m-1}\Omega_{g,m}$ is an exact one-form (even for every $m$). The residue of an exact one-form vanishes at each of its poles, that gives
@@
We can subtract this from (2.11) multiplied by $y(p)=\log(p)$ for any $p\in\{1,\frac{1}{\alpha_{j}},\frac{1}{\beta_{j}}\}$
@@
where the integration constant is irrelevant. On the RHS, we have several terms, where each of them is not an exact one-form, but the whole RHS is still exact. However, due to the carefully chosen constant $y(p)$, we have that (y(q)y(p))(y(q)-y(p)) vanishes linearly in (2.12). Thus, the first term on the RHS of (2.13) does not contribute, also $\omega_{g-1,2}^{\vee}=0$ for $g>0$. Finally, we end up with
@@
Sending the Kähler parameters to zero just the first term survives with the prefactor $1+r+s$ corresponding to the Euler characteristic of the toric Calabi-Yau. Thus, this computation gives an independent proof for [26, Conjecture 3] for strip geometries.
@@
Interchanging the $\hbar$-sum and the sum over $n$, one can resum for any $n$ the sum over $g$ (formally),
$\frac{(\frac{\beta_{i}}{\beta_{j}})^{n}}{n}\sum_{g=1}^{\infty}\hbar^{2g-2}\frac{n^{2g-2}B_{2g}}{2g(2g-2)!}=\frac{(\frac{\beta_{i}}{\beta_{j}})^{n}}{n(e^{\frac{\hbar n}{2}}-e^{-\frac{\hbar n}{2}})^{2}}.$
@@
where the indices $(i,j)$ range from 0 to $s$ for $\beta$’s and $(i,j)$ range from 1 to $r$ for $\alpha$’s. This is the result in one sector. A proper mathematical derivation would be via resurgence applying Borel summation techniques from [41] and the final result would depend on the integration contour [11].
@@
In this section, we compare the explicit results obtained from TR in (2.15) with the free energy of topological string theory (1.3) via the identification (1.8) of 5D BPS index. Our proposal amounts to computing the 5D BPS indices for strip geometries using the machinery of TR and the newly established xxyy duality [3, 46, 48] together with Log-TR [47, 6].
Starting from $Z_{\mathrm{top}}$, following [74], one can compute the BPS partition function in the noncommutative chamber as
@@
for these geometries, where the BPS partition function was shown to agree with matrix model computations. More precisely, one should perform Borel resummation on the asymptotic series (2.15) in different sectors of the Borel plane following [42, 11]. This way, one lands in one of the DT chambers [78, 54, 74]. Writing $Z_{\mathrm{top}}$ as in (1.10), one can extract the DT invariants. In fact, these expressions are asymptotic in $\hbar$, but are convergent in $Q_{i}$. The convergence conditions on $Q_{i}$ determine the DT chamber, as was also observed in the computations of Nekrasov partition functions [71, 72]. In our expressions (2.15) and (2.16), we keep the parameters $\alpha_{i}$ and $\beta_{i}$ quite general, without specifying the exact combinations $Q_{i}$’s, the complex structure parameters that correspond to them. The right identification of the parameters is made once we consider the convergence of such series in $Q_{i}$’s.
@@
There are a few already established techniques for computing these 5D indices using other methods. One of them is the “exponential networks” [30, 15, 17, 18, 19]. Briefly, the idea is to study the 3D-5D coupled system that we introduced in §1. Then one analyzes the wall-crossing of the 3D-5D BPS states. The information of the 5D BPS spectrum can be extracted by constructing a nonabelianization map for the exponential networks. In other words, the wall-crossing formula for 3D-5D BPS states, which is of the universal Kontsevich-Soibelman form [61], one can extract 5D BPS indices, by looking at wall-crossing of some particular 3D-5D BPS states and this is achieved by analyzing the topologies of the exponential networks associated to the corresponding toric CY threefolds. There are several other techniques to compute these invariants, for example using the BPS quiver and studying the moduli space of stable representations therein [20, 68]. Moreover, there exist techniques for computing them using supersymmetric localization [71, 12] or using methods from matrix model [74, 77], gluing of topological vertex [50], to name a few. The strip geometries are rather well-studied and many of the techniques mentioned above can extract the 5D BPS indices quite efficiently. Each of these methods brings forth certain aspects of mirror symmetry, such as the counting in [17, 16] are related to A-branes which are objects in Fukaya category, whereas the BPS quivers in [68] are related to counting of B-branes which are objects in the derived category of coherent sheaves. This exemplifies one aspect of homological mirror symmetry.
@@
Advent of supersymmetric localization [72, 73, 55] gave rise to yet another set of powerful techniques to study these 5D indices. In this situation, one needs to study certain moduli space which are multiplicative Higgs bundles. This is a more direct path towards quantization and as we shall see in an upcoming work [13] topological recursion is useful in that context too. Finally, there are several studies to understand the integrable structures of these 5D theories [70, 69, 74].
@@
The direct computation using TR has also its own implications. GW invariants can be reformulated using “remodeling conjecture” [25] on CY threefold $\mathcal{X}$. The result $F=\sum_{g\geq 0}\hbar^{2g-2}F_{g}$ is an asymptotic series in $\hbar$, where coefficients of each of the terms in $F_{g}$ in (2.15) encode the Gopakumar-Vafa invariants, for strip geometries. These GV invariants have clear interpretation in M-theory, as counting of M2-branes. The counts are also equal to the 5D BPS indices. In this work, we perform a verification of these statements through topological recursion. We provide some checks of this fact by comparing the free energy (2.15) and 5D invariants computed using other techniques, showing that they are in accordance with each other. [^1] The computations of the 5D indices referred to in the literature are carried out on both the A–model and B–model sides of these geometries: on the A–model (Fukaya category) side one relies on the techniques of exponential networks, while on the B–model (derived category of coherent sheaves) side one uses quiver descriptions and methods from representation theory. A thorough physical justification connecting both techniques is now available in the literature. In this sense, recovering these numbers directly from Topological Recursion can be interpreted as a “proof via TR”. One crucial simplification for the cases here is that, because there is no compact four cycle, the wall-crossing of the 5D BPS states is tractable and we can extract them by essentially analyzing one of the wall-crossing chambers.